Find the equation of the line that contains the given point and has the given slope. Let y be the dependent variable and let x be the independent variable.
Point (1, 1), m = 4/5

Answer by dabanfield(803)

Find the equation of the line that contains the given point and has the given slope. Let y be the dependent variable and let x be the independent variable.
Point (1, 1), m = 4/5
The standard form of a line in slope-intercept form is y = m*x + b where m is the slope and b is the value of y where the line crosses the x-axis (that is when x=0).
In this case we have slope = 4/5 so the line looks like:
y = (4/5)*x + b
We still need to find the value for b. Since the point (1,1) is a point on the line we know when x = 1, then y = 1. Substituting these values for x and y in the equation above we have:
1 = (4/5)*1 + b
1 = 4/5 + b
b = 1 - 4/5 = 1/5
So the equation is: y = (4/5)*x + 1/5